OpenStudy (kj4uts):

1. A flask contains 4.00 g of solid sodium hydroxide. Some water is added to the flask to dissolve the sodium hydroxide, and enough water is added to fill the flask to a volume of 50.0 mL. What volume of 1.00 M HCl is needed to neutralize the sodium hydroxide? A. 50.0 mL B. 100.0 mL C. 200.0 mL 2. A beaker contains 100.0 mL of a 0.100 M solution of sulfuric acid. When 100.0 mL water is added to this solution, what is the concentration of the diluted sulfuric acid? A. 0.0250 M B. 0.0500 M C. 0.200 M 3. How many moles of NaOH are required to react completely with 1.00 mole of nitric acid? A. 1.00 mol B. 0.500 mol C. 2.000 mol Please explain. Thank you!

OpenStudy (aaronq):

1. Neutralization implies that the moles of base are equal to the moles of acid, and thus they "cancel" each other by reacting to form water. \(OH^-+H^+\rightleftharpoons H_2O \) You need to find the moles of NaOH, these will be the moles of HCl you need. Then just find the volume of HCl solution needed.

OpenStudy (kj4uts):

@aaronq I found the moles of NaOH to be 39.997 g/mol

OpenStudy (kj4uts):

The molar mass of HCl is 36.46094 g/mol, I am not sure what to do for number 3.

OpenStudy (aaronq):

that's the molar mass not the moles

OpenStudy (aaronq):

\(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (kj4uts):

Ok so I have the molar mass, moles = mass / 39.997? How do I find the moles, is that the "1.00 mole"?

OpenStudy (aaronq):

So the moles of NaOH are 0.1 moles \(\sf moles =\dfrac{ 4 g}{39.997 g/mol}\)

OpenStudy (aaronq):

Now we know that they react in a 1:1 ratio (you would have to write the equation for the reaction and balance it). So the moles of NaOH = moles of HCl Now to find the volume of HCl

OpenStudy (kj4uts):

4/39.337=0.1000

OpenStudy (kj4uts):

Is there a formula I use to find the volume of HCI

OpenStudy (aaronq):

the molarity formula, since those are the units the concentration of the solution is in \(\sf Molarity=\dfrac{moles~of~solute}{L~of ~solution}\)

OpenStudy (aaronq):

gtg